OpenStudy (maeganr):

#14. Write an equation of a line that has the slope of -1/2 in each form. a. point-slope form b. slope-intercept form

OpenStudy (maeganr):

@erdog82 @Butterfly16 @chihiroasleaf

OpenStudy (anonymous):

First, I believe you have to decide what points you want to use to get the slope of -1/2

OpenStudy (maeganr):



OpenStudy (maeganr):

could this help me?

OpenStudy (anonymous):

some points could be (4,-1) (6,-2)

OpenStudy (maeganr):

so how do i write an equation of a line that has a slope of -1/2 in point-slope form, and slope-intercept form.

OpenStudy (anonymous):

well what is y1

OpenStudy (maeganr):

-1 ?

OpenStudy (anonymous):

yup, so so far we have y+1= -1/2 x - x1 what is x1?

OpenStudy (maeganr):

4

OpenStudy (maeganr):

?

OpenStudy (anonymous):

yup, so then we have y+1=-1/2 x -4

OpenStudy (maeganr):

thats it for a? we don't have to solve?! now how do we do b? y=mx+b

OpenStudy (anonymous):

what did you have to do for 13 on your last question?

OpenStudy (maeganr):

what do you mean?

OpenStudy (anonymous):

you have to subtract 1 from -4

OpenStudy (maeganr):

so its -5

OpenStudy (anonymous):

so then what do you get?

OpenStudy (maeganr):

-5 ?

OpenStudy (anonymous):

so then its y=-1/2 x - 5

OpenStudy (anonymous):

and your done!

OpenStudy (maeganr):

now i have 10 problems of standard form to do and 15 problems to do of parallel and perpendicular lines but i don't want to keep bothering you.

OpenStudy (anonymous):

oh, don't worry, it will help me! I am in "Enriched Algebra" and I just learned that, and I understand it really well - helping you will help me study and practice!